# Detecting Change Points of Covariance Matrices in High Dimensions

### v1

@@ 4.1 Proof of Theorem 1
$\log|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}|$ $=\log\big{|}\frac{n}{n-1}\hat{\mathbf{\Sigma}}\big{|}+\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\mathbf{\Sigma}}^{-1}\overline{\mathbf{y}}\right)$
$=-p\log\left(1-\frac{1}{n}\right)+\log\left|\hat{\mathbf{\Sigma}}\right|+\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\mathbf{\Sigma}}^{-1}\overline{\mathbf{y}}\right).$
A Taylor expansion shows that $-p\log\left(1-\frac{1}{n}\right)=p/n+o(1)$, and it also holds $\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\mathbf{\Sigma}}^{-1}\overline{\mathbf{y}}\right)=\log(1-p/n)+o(1)$ almost surely (see Section 4.3.1 in Heiny and Parolya, 2023). Thus, we obtain
$\log|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}|$ $=\log\left|\hat{\mathbf{\Sigma}}\right|+\frac{p}{n}+\log\left(1-\frac{p}{n}\right)+o(1)\quad\textnormal{almost surely}.$ (4.2)
@@
$\log|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}|$ $=\log\left|\hat{\mathbf{\Sigma}}_{1:\lfloor nt\rfloor}\right|+\frac{p}{\lfloor nt\rfloor}+\log\Big{(}1-\frac{p}{\lfloor nt\rfloor}\Big{)}+o(1)$ (4.3)
$\log|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}|$ $=\log\left|\hat{\mathbf{\Sigma}}_{(\lfloor nt\rfloor+1):n}\right|+\frac{p}{n-\lfloor nt\rfloor}+\log\Big{(}1-\frac{p}{n-\lfloor nt\rfloor}\Big{)}+o(1)$ (4.4)
@@
$\frac{2}{n}\log\Lambda_{n,t}^{\textnormal{cen}}$ $=\frac{\lfloor nt\rfloor}{n}\log\left|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}\right|+\frac{n-\lfloor nt\rfloor}{n}\log\left|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}\right|-\log\left|\hat{\mathbf{\Sigma}}^{\textnormal{cen}}\right|$
$=\frac{\lfloor nt\rfloor}{n}\log\left|\hat{\mathbf{\Sigma}}_{1:\lfloor nt\rfloor}\right|+\frac{n-\lfloor nt\rfloor}{n}\log\left|\hat{\mathbf{\Sigma}}_{(\lfloor nt\rfloor+1):n}\right|-\log\left|\hat{\mathbf{\Sigma}}\right|+\log\Big{(}1-\frac{p}{\lfloor nt\rfloor}\Big{)}$
@@ 4.1.1 Proof of Theorem 4 - weak convergence of finite-dimensional distributions
In the following, we prove Theorem 4, and the necessary auxiliary results are stated in Section 4.1.2.
@@
that is, if we let $\mathbf{X}_{i,j:k}=(\mathbf{b}_{1,j:k},\ldots,\mathbf{b}_{i,j:k})^{\top}_{\in\mathbb{R}^{k\times(k-j+1)\times i}},$ then
@@
Note that $X_{i:j}=X_{p:i,j}$, set $\mathbf{P}(0;j:k)=\mathbf{I}$ and $\mathbf{P}(i;1:n)=\mathbf{P}(i).$ Before rewriting $\log\Lambda_{n,t}$, we need some preparations. By applying QR-decompositions to $\mathbf{X}_{1:\lfloor nt\rfloor}^{\top},~{}\mathbf{X}_{(\lfloor nt\rfloor+1):n}^{\top}$ and $\mathbf{X}_{n}^{\top}$ (see (Wang, Han and Pan, 2018, Section 2) for more details), respectively, we have
$\left|n\hat{\mathbf{I}}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i}^{\top}\mathbf{P}(i-1)\mathbf{b}_{i},\quad$
$\left|\lfloor nt\rfloor\hat{\mathbf{I}}_{1:\lfloor nt\rfloor}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor},\quad$
$\left|(n-\lfloor nt\rfloor)\hat{\mathbf{I}}_{(\lfloor nt\rfloor+1):n}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}.$
Thus, under the null hypothesis of no change point, the likelihood ratio statistic does not depend on $\mathbf{\Sigma}$ and we may write
@@
$=\lfloor nt\rfloor\log\left|\hat{\mathbf{I}}_{1:\lfloor nt\rfloor}\right|+(n-\lfloor nt\rfloor)\log\left|\hat{\mathbf{I}}_{(\lfloor nt\rfloor+1):n}\right|-n\log\left|\hat{\mathbf{I}}\right|$
$=\lfloor nt\rfloor\log\left|\lfloor nt\rfloor\hat{\mathbf{I}}_{1:\lfloor nt\rfloor}\right|+(n-\lfloor nt\rfloor)\log\left|(n-\lfloor nt\rfloor)\hat{\mathbf{I}}_{(\lfloor nt\rfloor+1):n}\right|-n\log\left|n\hat{\mathbf{I}}\right|$ (4.10)
@@
$=\lfloor nt\rfloor\sum_{i=1}^{\lfloor nt\rfloor}\log\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor}$
$+(n-\lfloor nt\rfloor)\sum_{i=\lfloor nt\rfloor+1}^{n}\log\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}$
$-n\sum_{i=1}^{n}\log\mathbf{b}_{i}^{\top}\mathbf{P}(i-1)\mathbf{b}_{i}+np\log n-\lfloor nt\rfloor p\log\lfloor nt\rfloor-(n-\lfloor nt\rfloor)p\log(n-\lfloor nt\rfloor).$ (4.11)
@@
$\log n!=n\log n-n+\frac{1}{2}\log(2\pi n)+\frac{1}{12n}+\mathcal{O}\left(n^{-3}\right),\quad n\to\infty,$
@@
where we applied Lemma 2 and Lemma 1 for the last estimate, which are given in Section 4.1.2 below. Defining
@@
$=\sum_{i=1}^{p}\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{(k_{1}-j_{1}-i+1)(k_{2}-j_{2}-i+1)}\mathbb{E}\Big{[}\prod_{l=1,2}\left\{\mathbf{b}_{i,j_{l}:k_{l}}^{\top}\mathbf{P}(i-1;j_{l}:k_{l})\mathbf{b}_{i,j_{l}:k_{l}}-(k_{l}-j_{l}-i+1)\right\}\big{|}\mathcal{A}_{i-1}\Big{]}$
@@
×$\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{1}:k_{1})\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{2}:k_{2})\right),$
@@
×$\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{1}:k_{1})\odot\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{2}:k_{2})\right)$
@@
For later use, we note that the $o(1)$ term in (4.24) does not depend on $t\in[t_{0},1-t_{0}]$, if we set $j_{1}=j_{2}=1,k_{1}=k_{2}=\lfloor nt\rfloor$ or $j_{1}=j_{2}=\lfloor nt\rfloor+1,k_{1}=k_{2}=n.$ Moreover, in the case $k_{2}-j_{2}+1=(k_{1}-j_{1}+1)\vee(k_{2}-j_{2}+1)$, it follows from Lemma 3 in Section 4.1.2 below that
@@
To calculate $\operatorname{cov}(Z(t_{1}),Z(t_{2}))$ using (4.24), we use that $\sigma^{2}(j_{1},k_{1},j_{2},k_{2})=0$ if $1\leq j_{1}\leq k_{1}<j_{2}\leq k_{2}\leq n$ (this corresponds to the case that $X_{i,j_{1}:k_{1}}$ and $X_{i,j_{2}:k_{2}}$ are independent and thus $\mathbb{E}[X_{i,j_{1}:k_{1}}X_{i,j_{2}:k_{2}}|\mathcal{A}_{i-1}]=0$ for all $1\leq i\leq p$. ). In the following, we assume that $t_{1}<t_{2}$, which implies $\sigma^{2}(1,\lfloor nt_{1}\rfloor,\lfloor nt_{2}\rfloor+1,n)=0$. Combining (4.23) and (4.25) gives
@@
Here, we used (4.25) to see that the contributions of the $\sigma_{2}^{2}$-terms cancel each other out. Next, we use Lemma 4 in Section 4.1.2 below to compute the term $\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)$. For all remaining $\sigma_{1}^{2}$-terms, we use (4.24) and obtain
@@
4.1.2 Auxiliary results for the proof of Theorem 4
The convergence of the finite-dimensional distributions is facilitated by the following auxiliary results, whose proofs are postponed to the supplementary material. To begin with, we have an result on the quadratic term appearing in the expansion of the test statistic.
@@ 4.1.2 Auxiliary results for the proof of Theorem 4
The following result shows that the logarithmic terms are negligible at a $\delta$-dependent rate. It will also be used in Section 4.1.3 when the proving the asymptotic tightness given in Theorem 5.
@@
Assume that (A1) and (A2) with some $\delta>0$ are satisfied. Then, it holds for all $t\in[t_{0},1-t_{0}]$
@@
4.1.3 Proof of Theorem 5 - asymptotic tightness of $(\mathbf{\Lambda}_{n})_{n\in\mathbb{N}}$
@@
Let Assumption (A1) and (A2) with some $\delta>0$ be satisfied and let $t_{1},t_{2}\in[t_{0},1-t_{0}]$ and $D_{i,j}$ be defined as in (4.18) for $j\in\{1,2\},~{}1\leq i\leq p.$ Then, there exists random variables $Z_{1}=Z_{1,n}(t_{1},t_{2}),Z_{2}=Z_{2,n}(t_{1},t_{2})$ such that
@@
If Assumption (A1) and (A2) with some $\delta>4$ are satisfied, then there exist random variables $Q_{n,1,t}$ and $Q_{n,2,t}$ with
$Q_{n,1,t}+Q_{n,2,t}=\sum_{i=1}^{p}\frac{\lfloor nt\rfloor}{n}\frac{X_{i,1:\lfloor nt\rfloor}}{2}^{2}+\frac{n-\lfloor nt\rfloor}{n}\sum_{i=1}^{p}\frac{X_{i,(\lfloor nt\rfloor+1):n}^{2}}{2}-\sum_{i=1}^{p}\frac{X_{i}^{2}}{2}-\frac{\breve{\sigma}_{n,t}^{2}}{2},$ (4.26)
@@
$\sup_{t\in[t_{0},1-t_{0}]}\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\lesssim\frac{1}{n^{1+\delta/8}}.$ (4.27)
Finally, we recall Lemma 2 given in Section 4.1.2 on the logarithmic terms. Using these auxiliary results, we are in the position to give a proof of Theorem 5.
@@
Note that under the moment assumption (A2) with $\delta>4$, we have by Lemma 2 and Lemma 6
$\sup_{t\in[t_{0},1-t_{0}]}\left(\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\vee\mathbb{E}|R_{n}(t)|\right)\lesssim\frac{1}{n^{1+d}}$ (4.28)
for some $d>0$, which may be chosen such that it coincides with the $d>0$ from Lemma 5. Note that if $t-r<1/n$, we have $\lfloor nr\rfloor=\lfloor ns\rfloor$ or $\lfloor ns\rfloor=\lfloor nt\rfloor$, and thus, $m(r,s,t)=0$ almost surely. If $t-r\geq 1/n,$ it holds for all $\lambda>0$ by Lemma 5 and (4.28),
@@
$\mathbb{P}(|L_{n,t}-L_{n,s}|>\lambda)\lesssim\left(t-s+\frac{1}{n}\right)^{1+d}+\frac{1}{n^{1+d}}.$ (4.30)
@@
Combining (4.29) and (4.30) with Corollary A.4 in Dette and Tomecki (2019), we have for $\lfloor mt_{0}\rfloor\leq j\leq\lfloor m(1-t_{0})\rfloor$
$\mathbb{P}\left(\sup_{t_{1},t_{2}\in K_{j}}\left|L_{n,t_{1}}-L_{n,t_{2}}\right|>\lambda\right)\lesssim\frac{1}{m^{1+d}}+\left(\frac{1}{n}+\frac{1}{m}\right)^{1+d}+\frac{1}{n^{1+d}}.$ (4.31)

### v2

@@ 4.1 Proof of Theorem 1
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}|$ $=\log\big|\frac{n}{n-1}\hat{\boldsymbol{\Sigma}}\big|+\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\boldsymbol{\Sigma}}^{-1}\overline{\mathbf{y}}\right)$
$=-p\log\left(1-\frac{1}{n}\right)+\log\left|\hat{\boldsymbol{\Sigma}}\right|+\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\boldsymbol{\Sigma}}^{-1}\overline{\mathbf{y}}\right).$
A Taylor expansion shows that $p\log\left(1-\frac{1}{n}\right)=p/n+o(1)$, and it also holds $\log\left(1-\overline{\mathbf{y}}^{\top}\hat{\boldsymbol{\Sigma}}^{-1}\overline{\mathbf{y}}\right)=\log(1-p/n)+o(1)$ almost surely (see Section 4.3.1 in Heiny and Parolya, 2024). Thus, we obtain
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}|$ $=\log\left|\hat{\boldsymbol{\Sigma}}\right|+\frac{p}{n}+\log\left(1-\frac{p}{n}\right)+o(1)\quad\textnormal{almost surely}.$ (4.2)
@@
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}|$ $=\log\left|\hat{\boldsymbol{\Sigma}}_{1:\lfloor nt\rfloor}\right|+\frac{p}{\lfloor nt\rfloor}+\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)+o(1)$ (4.3)
$\log|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}|$ $=\log\left|\hat{\boldsymbol{\Sigma}}_{(\lfloor nt\rfloor+1):n}\right|+\frac{p}{n-\lfloor nt\rfloor}+\log\Big(1-\frac{p}{n-\lfloor nt\rfloor}\Big)+o(1)$ (4.4)
@@
$\frac{2}{n}\log\Lambda_{n,t}^{\textnormal{cen}}$ $=\frac{\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{1:\lfloor nt\rfloor}\right|+\frac{n-\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}_{(\lfloor nt\rfloor+1):n}\right|-\log\left|\hat{\boldsymbol{\Sigma}}^{\textnormal{cen}}\right|$
$=\frac{\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}_{1:\lfloor nt\rfloor}\right|+\frac{n-\lfloor nt\rfloor}{n}\log\left|\hat{\boldsymbol{\Sigma}}_{(\lfloor nt\rfloor+1):n}\right|-\log\left|\hat{\boldsymbol{\Sigma}}\right|+\log\Big(1-\frac{p}{\lfloor nt\rfloor}\Big)$
@@ 4.1.1 Proof of Theorem 2 - weak convergence of finite-dimensional distributions
In the following, we prove Theorem 2, and the necessary auxiliary results are stated in Section A.1.
@@
that is, if we let $\mathbf{X}_{i,j:k}=(\mathbf{b}_{1,j:k},\ldots,\mathbf{b}_{i,j:k})^{\top}_{\in\mathbb{R}^{i\times(k-j+1)}},$ then
@@
Note that $X_{i:j}=X_{p:i,j}$, set $\mathbf{P}(0;j:k)=\mathbf{I}$ and $\mathbf{P}(i;1:n)=\mathbf{P}(i).$ Before rewriting $\log\Lambda_{n,t}$, we need some preparations. By applying QR-decompositions to $\mathbf{X}_{1:\lfloor nt\rfloor}^{\top},~\mathbf{X}_{(\lfloor nt\rfloor+1):n}^{\top}$ and $\mathbf{X}_{n}^{\top}$ (see (Wang et al., 2018, Section 2) for more details), respectively, we have
[…]
$\left|\lfloor nt\rfloor\hat{\mathbf{I}}_{1:\lfloor nt\rfloor}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor},\quad$ (4.10)
$\left|(n-\lfloor nt\rfloor)\hat{\mathbf{I}}_{(\lfloor nt\rfloor+1):n}\right|$ $=\prod_{i=1}^{p}\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}.$ (4.11)
Thus, under the null hypothesis of no change point, the likelihood ratio statistic does not depend on $\boldsymbol{\Sigma}$ and we may write
@@
[…]
@@
$=\lfloor nt\rfloor\sum_{i=1}^{p}\log\mathbf{b}_{i,1:\lfloor nt\rfloor}^{\top}\mathbf{P}(i-1;1:\lfloor nt\rfloor)\mathbf{b}_{i,1:\lfloor nt\rfloor}$
$+(n-\lfloor nt\rfloor)\sum_{i=1}^{p}\log\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}^{\top}\mathbf{P}(i-1;(\lfloor nt\rfloor+1):n)\mathbf{b}_{i,(\lfloor nt\rfloor+1):n}$
$-n\sum_{i=1}^{p}\log\mathbf{b}_{i}^{\top}\mathbf{P}(i-1)\mathbf{b}_{i}+np\log n-\lfloor nt\rfloor p\log\lfloor nt\rfloor-(n-\lfloor nt\rfloor)p\log(n-\lfloor nt\rfloor).$ (4.13)
@@
$\log n!=n\log n-n+\frac{1}{2}\log(2\pi n)+\frac{1}{12n}+\mathcal{O}\left(n^{-3}\right),\quad n\to\infty,$ (4.18)
@@
where we applied Lemma 2 and Lemma 1 for the last estimate, which are given in Section A.1. Defining
@@
$=\sum_{i=1}^{p}\frac{(k_{1}-j_{1}+1)(k_{2}-j_{2}+1)}{(k_{1}-j_{1}-i+1)(k_{2}-j_{2}-i+1)}$ (4.26)
$\quad\quad\times\mathbb{E}\Big[\prod_{l=1,2}\left\{\mathbf{b}_{i,j_{l}:k_{l}}^{\top}\mathbf{P}(i-1;j_{l}:k_{l})\mathbf{b}_{i,j_{l}:k_{l}}-(k_{l}-j_{l}-i+1)\right\}\big|\mathcal{A}_{i-1}\Big]$
@@
$\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{1}:k_{1})\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{2}:k_{2})\right),$
@@
$\quad\quad\times\operatorname{tr}\left(\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{1}:k_{1})\odot\mathbf{P}^{(j_{1}\vee j_{2}):(k_{1}\wedge k_{2})}(i-1;j_{2}:k_{2})\right)$
@@
For later use, we note that the $o(1)$ term in (4.28) does not depend on $t\in[t_{0},1-t_{0}]$, if we set $j_{1}=j_{2}=1,k_{1}=k_{2}=\lfloor nt\rfloor$ or $j_{1}=j_{2}=\lfloor nt\rfloor+1,k_{1}=k_{2}=n.$ Moreover, in the case $k_{2}-j_{2}+1=(k_{1}-j_{1}+1)\vee(k_{2}-j_{2}+1)$, it follows from Lemma 3 in Section A.1 below that
@@
To calculate $\operatorname{cov}(Z(t_{1}),Z(t_{2}))$ using (4.28), we use that $\sigma^{2}(j_{1},k_{1},j_{2},k_{2})=0$ if $1\leq j_{1}\leq k_{1}<j_{2}\leq k_{2}\leq n$ (this corresponds to the case that $X_{i,j_{1}:k_{1}}$ and $X_{i,j_{2}:k_{2}}$ are independent and thus, for all $1\leq i\leq p$, $\mathbb{E}[X_{i,j_{1}:k_{1}}X_{i,j_{2}:k_{2}}|\mathcal{A}_{i-1}]=0$). In the following, we assume that $t_{1}<t_{2}$, which implies $\sigma^{2}(1,\lfloor nt_{1}\rfloor,\lfloor nt_{2}\rfloor+1,n)=0$. Combining (4.27) and (4.29) gives
@@
Here, we used (4.29) to see that the contributions of the $\sigma_{2}^{2}$-terms cancel each other out. Next, we use Lemma 4 in Section A.1 below to compute the term $\sigma_{1}^{2}(\lfloor nt_{1}\rfloor+1,n,1,\lfloor nt_{2}\rfloor)$. For all remaining $\sigma_{1}^{2}$-terms, we use (4.28) and obtain
@@
{funding}
This work was partially supported by the DFG Research unit 5381 Mathematical Statistics in the Information Age, project number 460867398, and the Aarhus University Research Foundation (AUFF), project numbers 47221 and 47388.
[…]
The online supplement contains the proofs of Theorem 1 and its auxiliary results.
A Supplementary Material
A.1 Auxiliary results for the proof of Theorem 2
The convergence of the finite-dimensional distributions is facilitated by the following auxiliary results, whose proofs are postponed to Section A.3. To begin with, we have a result on the quadratic term appearing in the expansion of the test statistic.
@@ A.1 Auxiliary results for the proof of Theorem 2
The following result shows that the logarithmic terms are negligible at a $\delta$-dependent rate. It will also be used in Section A.2 when the proving the asymptotic tightness given in Theorem 3.
@@
Assume that (A-1) and (A-2) with some $\delta>0$ are satisfied. Then, it holds for all $t\in[t_{0},1-t_{0}]$
@@
A.2 Proof of Theorem 3 - asymptotic tightness
@@
Let Assumption (A-1) and (A-2) with some $\delta>0$ be satisfied and let $t_{1},t_{2}\in[t_{0},1-t_{0}]$ and $D_{i,j}$ be defined as in (4.21) for $j\in\{1,2\},~1\leq i\leq p.$ Then, there exists random variables $Z_{1}=Z_{1,n}(t_{1},t_{2}),Z_{2}=Z_{2,n}(t_{1},t_{2})$ such that
@@
If Assumption (A-1) and (A-2) with some $\delta>4$ are satisfied, then there exist random variables $Q_{n,1,t}$ and $Q_{n,2,t}$ with
$Q_{n,1,t}+Q_{n,2,t}=\sum_{i=1}^{p}\frac{\lfloor nt\rfloor}{n}\frac{X_{i,1:\lfloor nt\rfloor}}{2}^{2}+\frac{n-\lfloor nt\rfloor}{n}\sum_{i=1}^{p}\frac{X_{i,(\lfloor nt\rfloor+1):n}^{2}}{2}-\sum_{i=1}^{p}\frac{X_{i}^{2}}{2}-\frac{\breve{\sigma}_{n,t}^{2}}{2},$ (A.1)
@@
$\sup_{t\in[t_{0},1-t_{0}]}\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\lesssim\frac{1}{n^{1+\delta/8}}.$ (A.2)
Finally, we recall Lemma 2 given in Section A.1 on the logarithmic terms. Using these auxiliary results, we are in the position to give a proof of Theorem 3.
@@
Note that under the moment assumption (A-2) with $\delta>4$, we have by Lemma 2 and Lemma 6
$\sup_{t\in[t_{0},1-t_{0}]}\left(\mathbb{E}|Q_{2,n,t}|^{2+\delta/4}\vee\mathbb{E}|R_{n}(t)|\right)\lesssim\frac{1}{n^{1+d}}$ (A.3)
for some $d>0$, which may be chosen such that it coincides with the $d>0$ from Lemma 5. Note that if tr<1/nt-r<1/n, we have $\lfloor nr\rfloor=\lfloor ns\rfloor$ or $\lfloor ns\rfloor=\lfloor nt\rfloor$, and thus, $m(r,s,t)=0$ almost surely. If $t-r\geq 1/n,$ it holds for all $\lambda>0$ by Lemma 5 and (A.3),
@@
$\mathbb{P}(|L_{n,t}-L_{n,s}|>\lambda)\lesssim\left(t-s+\frac{1}{n}\right)^{1+d}+\frac{1}{n^{1+d}}.$ (A.5)
@@
Combining (A.4) and (A.5) with Corollary A.4 in Dette and Tomecki (2019), we have for $\lfloor mt_{0}\rfloor\leq j\leq\lfloor m(1-t_{0})\rfloor$
$\mathbb{P}\left(\sup_{t_{1},t_{2}\in K_{j}}\left|L_{n,t_{1}}-L_{n,t_{2}}\right|>\lambda\right)\lesssim\frac{1}{m^{1+d}}+\left(\frac{1}{n}+\frac{1}{m}\right)^{1+d}+\frac{1}{n^{1+d}}.$ (A.6)